# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ 4 Numerical Experiments
In this section, we conduct numerical experiments on both a two-dimensional triangular lattice and a three-dimensional body-centered cubic (BCC) lattice. We introduce quadrupole edge dislocations into the lattice structure with periodic boundary conditions and compute the corresponding variables (e.g., $A^{2}$ defined in (2.16)) and the strain field. To further validate the proposed numerical methods, we compare the results with those obtained from continuum elasticity theory of dislocations [47, 48].
@@ 4.1 Two-dimensional triangular lattice
We first consider a two-dimensional triangular lattice described in Example 2.3. Let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By minimizing the energy functional given in (2.15), we obtain $\phi_{0}$ as follows:
$\phi_{0}=\frac{\tau+\sqrt{\tau^{2}-15\nu B_{0}}}{15v}.$
Further computational details can be found in Appendix B.2. A square computational domain $L_{x}\times L_{y}$, with edge lengths $L_{x}=L_{y}=80\pi$, is utilized for numerical simulations. To simulate crystal dislocations, only horizontal shear stress parallel to the $x$-axis is applied, with no force exerted in the $y$-direction. The initial condition $\eta_{j,0}$ of the amplitude functions can be determined based on $\eta_{j}=\phi_{0}e^{\imath\bm{k_{j}}\cdot u(\bm{r})}$ by defining $u(\bm{r})$ as the displacement of the lattice. Given $\bm{r}:=x\widehat{\bm{x}}+y\widehat{\bm{y}}$ as the spatial position vector, the displacement $u(\bm{r})$ is given by:
$u(\bm{r})=\left\{\begin{aligned} &-u_{x}\widehat{\bm{x}},\quad&\frac{L_{y}}{4}<y<\frac{3L_{y}}{4}\\ &+u_{x}\widehat{\bm{x}},\quad&\text{elsewhere}\end{aligned}\right.,$ (4.1)
where $u_{x}=\frac{a_{x}}{L_{x}}x$ with $a_{x}=\frac{4\pi}{\sqrt{3}}$ and $u_{y}=0$. With this choice, a strain of $\varepsilon=\pm\frac{a_{x}}{L_{x}}$ is applied to the corresponding region.
We solve the APFC model (2.13) using pseudospectral method, to demonstrate the efficiency of the method, we present the CPU time (in seconds) required to solve the model for various values of $N$, which represent the degrees of freedom in each direction, as outlined in Section 3.1. The time step size, $\Delta t$, is set to 0.1. Table 1 shows the time required for different values of $N$. The CPU time scales approximately as $O(N^{1.16})$, demonstrating the high efficiency of our model solution process. This efficiency is further underscored by the relatively low computational effort required, as discussed in Section 3.2.
N 32 64 128 256 512 1024
@@ 4.1.1 Density $n$ and Amplitude Parameter $A^{2}$
By solving APFC model (2.13), we can obtain the amplitude functions $\eta_{j}$. These functions can then be used to reconstruct density $n$ by (2.2) and calculate $A^{2}$ by (2.16). Figure 2 presents the visualization of the density $n$ defined in (2.2) and the parameter $A^{2}$ defined in (2.16), using the parameters $B_{0}=0.02$, $B_{x}=0.98$, $v=\frac{1}{3}$, and $\tau=\frac{1}{3}$ from (2.15) as provided in [24]. Figure 2a and Figure 2b display the density $n$ of the perfect lattice and the defect lattice, which can be reconstructed from (2.2). In Figure 2b, the area near the defect is extracted, and the burgers vector can be clearly seen. Figure 2c depicts four blue ellipses representing the regions of defects formed as the system stabilizes, indicating the distribution of dislocations, specifically a quadrupole dislocation. This numerical distribution satisfies the property of the parameter $A^{2}$: it remains constant within ordered phases but decreases in the presence of defects.
@@
Figure 2: Triangular lattice: The illustration of the density $n$ defined in  (2.2) of the perfect lattice and defect lattice, and the parameter $A^{2}$ defined in  (2.16). Both are colored by values.
@@ 4.1.2 Strain Field and Comparison with Continuum Elasticity
$\bm{\varepsilon}=\frac{1}{2}\left[(\nabla u)+(\nabla u)^{T}\right].$ (4.2)
By solving the APFC model (2.13), we can obtain $\eta_{j}$. The relationship between $\eta_{j}$ and $u$ is established by introducing an intermediate variable:
$\varphi_{j}=\bm{k_{j}}\cdot u(\bm{r}),\quad j=1,\ldots,N.$ (4.3)
Thus, the equation (3.3) can be written as $\eta_{j}=\phi_{0}e^{\imath\varphi_{j}}$. Furthermore, $\varphi_{j}$ can be expressed in terms of $\eta_{j}$ as follows:
$\varphi_{j}=\arg(\eta_{j})=\arctan\left(\frac{\mathrm{Im}(\eta_{j})}{\mathrm{Re}(\eta_{j})}\right),$ (4.4)
where (4.3) defines a system of $N$ linear equations concerning $u$. We employ the least squares method to solve this system. After some straightforward calculations, the least squares solution of (4.3) in the sense of the Moore-Penrose inverse is expressed as follows,
$u(\bm{r})=\left\{\begin{array}[]{lr}u_{x}=-\frac{\sqrt{3}}{3}\varphi_{1}+\frac{\sqrt{3}}{3}\varphi_{3},\\ u_{y}=-\frac{1}{3}\varphi_{1}+\frac{2}{3}\varphi_{2}-\frac{1}{3}\varphi_{3}.\end{array}\right.$ (4.5)
To calculate the strain field according to (4.2), we need to obtain $\nabla u$. Therefore, based on the expression in (4.5), we compute the partial derivatives of $\varphi_{j}$,
$\frac{\partial\varphi_{j}}{\partial\xi}=\frac{1}{|\eta_{j}|^{2}}\left[\frac{\partial\mathrm{Im}(\eta_{j})}{\partial\xi}\mathrm{Re}(\eta_{j})-\frac{\partial\mathrm{Re}(\eta_{j})}{\partial\xi}\mathrm{Im}(\eta_{j})\right],\quad\xi=x,y.$ (4.6)
Thus, the strain field can be computed solely from $\eta_{j}$, obtained by solving the APFC model (2.13), without the need for additional data. The formulas for the components of the strain field are:
$\begin{split}\varepsilon_{xx}&=-\frac{\sqrt{3}}{3}\frac{\partial\varphi_{1}}{\partial x}+\frac{\sqrt{3}}{3}\frac{\partial\varphi_{3}}{\partial x},\\ \varepsilon_{yy}&=-\frac{1}{3}\frac{\partial\varphi_{1}}{\partial y}+\frac{2}{3}\frac{\partial\varphi_{2}}{\partial y}-\frac{1}{3}\frac{\partial\varphi_{3}}{\partial y},\\ \varepsilon_{xy}&=-\frac{\sqrt{3}}{3}\frac{\partial\varphi_{1}}{\partial y}+\frac{\sqrt{3}}{3}\frac{\partial\varphi_{3}}{\partial y}-\frac{1}{3}\frac{\partial\varphi_{1}}{\partial x}+\frac{2}{3}\frac{\partial\varphi_{2}}{\partial x}-\frac{1}{3}\frac{\partial\varphi_{3}}{\partial x}.\end{split}$ (4.7)
Figure 3c illustrates the components of the strain field in a two-dimensional triangular lattice. The distribution of the strain field for a quadrupole dislocation is observed, and notably, the field distribution at each defect core aligns with theoretical predictions discussed in prior work [39].
@@
Figure 3: The components of strain field for two-dimensional triangular lattice: (a) $\varepsilon_{xx}$, (b) $\varepsilon_{xy}$ and (c) $\varepsilon_{yy}$.
Comparison with Continuum Elasticity Theory
We compare the APFC solution with continuum elasticity theory to describe the elastic field of a dislocation [39], assuming that the dislocation characteristics, such as the Burgers vector $\bm{b}$ and the elastic constants, are known [49]. Without loss of generality, we set the position of the dislocation core at $\bm{x_{c}}=(0,0)$ and that the Burgers vector $\bm{b}$ aligns with the $x$-axis. To mitigate the singularity inherent at the dislocation core, we adopt the regularization method proposed in [47]. The stress field components can be computed as follows,
$\begin{split}\sigma_{xx}/\sigma_{0}&=-y(3\zeta^{2}+3x^{2}+y^{2}),\\ \sigma_{yy}/\sigma_{0}&=x(\zeta^{2}+x^{2}-y^{2}),\\ \sigma_{yx}/\sigma_{0}&=\sigma_{xy}/\sigma_{0}=x(\zeta^{2}+x^{2}-y^{2}),\end{split}$ (4.8)
@@
$\sigma_{0}=\frac{E|\bm{b}|}{2\pi(1-\nu^{2})(\zeta^{2}+x^{2}+y^{2})^{2}},$ (4.9)
with $\zeta=|\bm{b}|/2$ representing the regularization term, $|\bm{b}|=\frac{4\pi}{\sqrt{3}}$, and $E$ and $\nu$ representing Young’s modulus and Poisson’s ratio, respectively.
@@
$\bm{\varepsilon}^{\rm CE}=\left[\frac{1+\nu}{E}\right]\bm{\sigma}^{\rm tot}-\left(\frac{\nu}{E}\right){\rm tr}(\bm{\sigma}^{\rm tot})\bm{I},$ (4.10)
where $\sigma^{\rm tot}$ is the total stress field. The strain field in each direction is obtained as follows:
$\begin{split}\varepsilon_{xx}&=\frac{1}{E}(\sigma_{xx}-\nu\sigma_{yy}),\\ \varepsilon_{yy}&=\frac{1}{E}(\sigma_{yy}-\nu\sigma_{xx}),\\ \varepsilon_{xy}&=\frac{1+\nu}{E}\sigma_{xy}.\end{split}$ (4.11)
@@
As shown in Figure 5, the strain field from the APFC model (2.13) is in good agreement with those obtained from continuum elasticity. Notably, we observe variations around the dislocation cores along the $l_{1}$ and $l_{3}$ directions, which reflects the expected strain field behavior in the presence of dislocations. This demonstrates the effectiveness of the Fourier pseudospectral method in solving the APFC model (2.13), yielding results that closely align with elasticity theory.
@@
Figure 5: Triangular lattice: Strain fields compared with CE along different lines. The red solid lines represent the results from CE theory, while the black circles represent the results from the APFC model (2.13).
Far Field Decay
To verify the accuracy of the Fourier pseudospectral method for solving the APFC model (2.13), we consider a larger computational domain with a side length of $640\pi$. We focus on the case of $\varepsilon_{xx}$ in the direction $l_{1}$, noting that the strain field in other directions yields similar results. We demonstrate the results by comparing the decay of $\varepsilon_{xx}$ as a function of the distance $r$ from the dislocation core. As shown in Figure 6, both $\varepsilon_{xx}(\text{APFC})$ and $\varepsilon_{xx}(\text{CE})$, representing the results obtained using the APFC model (2.13) and continuum elasticity (CE) theory, respectively, decay like $r^{-1}$. The error between them decays as $r^{-2}$, which is consistent with the results shown in [50].
Figure 6: Triangular lattice: The decay of the strain field $\varepsilon_{xx}$ along the $l_{1}$ direction as a function of the distance $r$ from the dislocation core. The red line represents the results obtained from CE theory, while the green line shows the results from the APFC model (2.13). The blue line represents the error between them.
@@ 4.2 Three-dimensional BCC lattice
For the three-dimensional BCC lattice described in Example 2.3 , we first let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By minimizing the energy functional given in (2.15), we obtain $\phi_{0}$ as follows:
$\phi_{0}=\frac{2\tau+\sqrt{4\tau^{2}-45vB_{0}}}{45v}.$
@@
A square cube area $L_{x}\times L_{y}\times L_{z}$, with edge length $L_{x}=L_{y}=L_{z}=60\pi$, is utilized for numerical simulation. To simulate crystal dislocation, shear stress is applied as in the two-dimensional case. Only horizontal shear stress parallel to the $xy$ plane is applied, with no force exerted in the z direction. Based on $\eta_{j}=\phi_{0}e^{\imath\bm{k_{j}}\cdot u(\bm{r})}$, the initial condition $\eta_{j,0}$ of the amplitude functions can be determined by defining $u(\bm{r})$ as a displacement to the lattice. Given $\bm{r}:=x\widehat{\bm{x}}+y\widehat{\bm{y}}+z\widehat{\bm{z}}$ representting the spatial position vector, the definition of $u(\bm{r})$ is then given by:
$u(\bm{r})=\left\{\begin{aligned} -u_{x}\widehat{\bm{x}}-u_{y}\widehat{\bm{y}}&,&\frac{L_{z}}{4}<z<\frac{3L_{z}}{4}\\ +u_{x}\widehat{\bm{x}}+u_{x}\widehat{\bm{x}}&,&{\rm elsewhere}\end{aligned}\right.,$ (4.12)
where $u_{x}=a_{x}/L_{x}x$, $u_{y}=a_{y}/L_{y}y$ and $u_{z}=0$ with $a_{x}=a_{y}=2\pi\sqrt{2}$ and $L_{x}$, $L_{y}$ being the length of computational domain. With this choice, a strain of
$\varepsilon_{x}=\pm a_{x}/L_{x},\quad\varepsilon_{y}=\pm a_{y}/L_{y}.$ (4.13)

### v3

@@ 4 Numerical Experiments
In this section, we conduct numerical experiments on both a two-dimensional triangular lattice and a three-dimensional body-centered cubic (BCC) lattice. We introduce quadrupole edge dislocations into the lattice structure with periodic boundary conditions and compute the corresponding variables (e.g., $A^{2}$ defined in (16)) and the strain field. To further validate the proposed numerical methods, we compare the results with those obtained from continuum elasticity theory of dislocations Cai2006A ; olson2023elastic .
@@ 4.1 Two-dimensional triangular lattice
We first consider a two-dimensional triangular lattice described in Example 1. Let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By minimizing the energy functional given in (15), we obtain $\phi_{0}$ as follows:
$\phi_{0}=\frac{\gamma+\sqrt{\gamma^{2}-15vB_{0}}}{15v}.$
Further computational details can be found in Appendix B.2. A square computational domain $L_{x}\times L_{y}$, with edge lengths $L_{x}=L_{y}=80\pi$, is utilized for numerical simulations. To simulate crystal dislocations, only horizontal shear stress parallel to the $x$-axis is applied, with no force exerted in the $y$-direction. The initial condition $\eta_{j,0}$ of the amplitude functions can be determined based on $\eta_{j}=\phi_{0}e^{\imath\bm{k}_{j}\cdot\bm{u}(\bm{r})}$ by defining $\bm{u}(\bm{r})$ as the displacement of the lattice. Given $\bm{r}:=x\widehat{\bm{x}}+y\widehat{\bm{y}}$ as the spatial position vector, the displacement $\bm{u}(\bm{r})$ is given by
$\bm{u}(\bm{r})=\left\{\begin{aligned} &-u_{x}\widehat{\bm{x}},\quad&\frac{L_{y}}{4}<y<\frac{3L_{y}}{4}\\ &+u_{x}\widehat{\bm{x}},\quad&\text{elsewhere}\end{aligned}\right$., (29)
where $u_{x}=(a_{x}/L_{x})x$ with $a_{x}=4\pi/\sqrt{3}$ and $u_{y}=0$. With this choice, a strain of $\varepsilon=\pm a_{x}/L_{x}$ is applied to the corresponding region.
We solve the APFC model (13) using the Fourier pseudospectral method. To validate the accuracy of the numerical scheme described in Section 3, a reference solution $\eta_{j}^{*}(\bm{r},T)$ is computed with a sufficiently small time step $\tau=1\times 10^{-4}$, a fine spatial mesh size $h=\pi/128$, and a final time $T=4$. The choice of $T$ is determined empirically so that the system attains a steady state, as confirmed by numerical experiments.
Let the reference solution of (13) be expressed as
$\eta_{j}^{*}(\bm{r},T)=\sum_{\bm{k}\in K_{N}^{d}}\eta_{j_{\bm{k}}}^{*}e^{\imath(B\bm{k})\cdot\bm{r}},\quad j=1,2,3,$ (30)
and define the numerical error by
$e_{j,N}=\|\eta_{j}^{*}(\cdot,T)-\tilde{\eta}_{j}(\cdot,T)\|_{L^{2}}^{2}=\sum_{\bm{k}\in K_{N}^{d}}\big|\eta_{j_{\bm{k}}}^{*}(T)-\tilde{\eta}_{j_{\bm{k}}}(T)\big|^{2}.$ (31)
Hence, the error of the Fourier pseudospectral method is measured by the convergence of the Fourier coefficients. Figure 2 shows the error $e_{j,N}$ for $j=1,2,3$ with $d=2$ as a function of $N$. For the nonlinear system under the assumptions of Theorem 2.1, $\eta_{j}$ is analytic, and the results confirm spectral accuracy, in agreement with Theorem 3.1.
Figure 2: Triangular lattice: numerical error $e_{j,N}$ for $j=1,2,3$, $d=2$ as a function of $N$.
To further demonstrate the computational efficiency, we report the CPU time (in seconds) required to solve the model for various values of $N$, which correspond to the degrees of freedom in each spatial direction (see Section 3.1). The time step size, $\Delta t$, is set to 0.1. Table 1 shows the time required for different values of $N$. The CPU time scales approximately as $O(N^{1.16})$, demonstrating the high efficiency of our model solution process. This efficiency is further underscored by the relatively low computational effort required, as discussed in Section 3.2.
Table 1: Triangular lattice: CPU time (s) vs. NN.
NN 32 64 128 256 512 1024
@@ 4.1.1 Density $n$ and amplitude parameter $A^{2}$
By solving APFC model (13), we can obtain the amplitude functions $\eta_{j}$. These functions can then be used to reconstruct density $n$ by (2) and calculate $A^{2}$ by (16). Figure 3 presents the visualization of the density $n$ defined in (2) and the parameter $A^{2}$ defined in (16), using the parameters $B_{0}=0.02$, $B_{x}=0.98$, $v=1/3$, and $\gamma=1/3$ from (15) as provided in 2017Controlling . Figure 3a and Figure 3b display the density $n$ of the perfect lattice and the defect lattice, which can be reconstructed from (2). In Figure 3b, the area near the defect is extracted, and the Burgers vector can be clearly seen. Figure 3c depicts four blue ellipses representing the regions of defects formed as the system stabilizes, indicating the distribution of dislocations, specifically a quadrupole dislocation. This numerical distribution satisfies the property of the parameter $A^{2}$: it remains constant within ordered phases but decreases in the presence of defects.
@@
Figure 3: Triangular lattice: The illustration of the density $n$ defined in (2) of the perfect lattice and defect lattice, and the parameter $A^{2}$ defined in (16). Both are colored by values.
@@ 4.1.2 Strain field and comparison with continuum elasticity
$\bm{\varepsilon}=\frac{1}{2}\left[(\nabla{\bm{u}})+(\nabla{\bm{u}})^{T}\right].$ (32)
By solving the APFC model (13), we can obtain $\eta_{j}$. The relationship between $\eta_{j}$ and $\bm{u}$} is established by introducing an intermediate variable:
$\varphi_{j}=\bm{k}_{j}\cdot\bm{u}(\bm{r}),\quad j=1,\ldots,M.$ (33)
Thus, the equation (21) can be written as $\eta_{j}=\phi_{0}e^{\imath\varphi_{j}}$. Furthermore, $\varphi_{j}$ can be expressed in terms of $\eta_{j}$ as follows:
$\varphi_{j}=\arg(\eta_{j})=\arctan\left(\frac{\mathrm{Im}(\eta_{j})}{\mathrm{Re}(\eta_{j})}\right),$ (34)
where (33) defines a system of $M$ linear equations related to $u$. We employ the least squares method to solve this system. After some straightforward calculations, the least squares solution of (33) in the sense of the Moore-Penrose inverse is expressed as follows:
$\bm{u}(\bm{r})=\left\{\begin{array}[]{lr}u_{x}=-\frac{\sqrt{3}}{3}\varphi_{1}+\frac{\sqrt{3}}{3}\varphi_{3},\\ u_{y}=-\frac{1}{3}\varphi_{1}+\frac{2}{3}\varphi_{2}-\frac{1}{3}\varphi_{3}.\end{array}\right$. (35)
To calculate the strain field according to (32), we need to obtain $\nabla{\bm{u}}$. To this end, based on the expression in (35), we compute the partial derivatives of $\varphi_{j}$,
$\frac{\partial\varphi_{j}}{\partial\xi}=\frac{1}{\lvert\eta_{j}\rvert^{2}}\left[\frac{\partial\mathrm{Im}(\eta_{j})}{\partial\xi}\mathrm{Re}(\eta_{j})-\frac{\partial\mathrm{Re}(\eta_{j})}{\partial\xi}\mathrm{Im}(\eta_{j})\right],\quad\xi=x,y.$ (36)
Thus, the strain field can be computed solely from $\eta_{j}$, obtained by solving the APFC model (13), without the need for additional data. The formulas for the components of the strain field are:
$\begin{split}\varepsilon_{xx}&=-\frac{\sqrt{3}}{3}\frac{\partial\varphi_{1}}{\partial x}+\frac{\sqrt{3}}{3}\frac{\partial\varphi_{3}}{\partial x},\\ \varepsilon_{yy}&=-\frac{1}{3}\frac{\partial\varphi_{1}}{\partial y}+\frac{2}{3}\frac{\partial\varphi_{2}}{\partial y}-\frac{1}{3}\frac{\partial\varphi_{3}}{\partial y},\\ \varepsilon_{xy}&=-\frac{\sqrt{3}}{3}\frac{\partial\varphi_{1}}{\partial y}+\frac{\sqrt{3}}{3}\frac{\partial\varphi_{3}}{\partial y}-\frac{1}{3}\frac{\partial\varphi_{1}}{\partial x}+\frac{2}{3}\frac{\partial\varphi_{2}}{\partial x}-\frac{1}{3}\frac{\partial\varphi_{3}}{\partial x}.\end{split}$ (37)
Figure 4 illustrates the components of the strain field in a two-dimensional triangular lattice. The distribution of the strain field for a quadrupole dislocation is observed, and notably, the field distribution at each defect core aligns with theoretical predictions discussed in prior work anderson2017theory .
@@
Figure 4: The components of strain field for two-dimensional triangular lattice: (a) $\varepsilon_{xx}$, (b) $\varepsilon_{xy}$, and (c) $\varepsilon_{yy}$.
Comparison with continuum elasticity theory.
We compare the APFC solution with continuum elasticity theory to describe the elastic field of a dislocation, assuming that the dislocation characteristics, such as the Burgers vector $\bm{b}$ and the elastic constants, are known anderson2017theory . Without loss of generality, we set the position of the dislocation core at $\bm{x}_{c}=(0,0)$ and that the Burgers vector $\bm{b}$ aligns with the $x$-axis. To mitigate the singularity inherent at the dislocation core, we adopt the regularization method proposed in Cai2006A . The stress field components can be computed as follows:
$\begin{split}\sigma_{xx}/\sigma_{0}&=-y(3\zeta^{2}+3x^{2}+y^{2}),\\ \sigma_{yy}/\sigma_{0}&=x(\zeta^{2}+x^{2}-y^{2}),\\ \sigma_{yx}/\sigma_{0}&=\sigma_{xy}/\sigma_{0}=x(\zeta^{2}+x^{2}-y^{2}),\end{split}$ (38)
@@
$\sigma_{0}=\frac{E\lvert\bm{b}\rvert}{2\pi(1-\nu^{2})(\zeta^{2}+x^{2}+y^{2})^{2}},$ (39)
with $\zeta=\lvert\bm{b}\rvert/2$ representing the regularization term, $\lvert\bm{b}\rvert=4/\sqrt{3}$, and $E$ and $\nu$ representing Young’s modulus and Poisson’s ratio, respectively.
@@
$\bm{\varepsilon}^{\rm CE}=\left[\frac{1+\nu}{E}\right]\bm{\sigma}^{\rm tot}-\left(\frac{\nu}{E}\right){\rm tr}(\bm{\sigma}^{\rm tot})\bm{I},$ (40)
where $\bm{\sigma}^{\rm tot}$ is the total stress field. The strain field in each direction is obtained as follows:
$\begin{split}\varepsilon_{xx}&=\frac{1}{E}(\sigma_{xx}-\nu\sigma_{yy}),\\ \varepsilon_{yy}&=\frac{1}{E}(\sigma_{yy}-\nu\sigma_{xx}),\\ \varepsilon_{xy}&=\frac{1+\nu}{E}\sigma_{xy}.\end{split}$ (41)
@@
As shown in Figure 6, the strain field from the APFC model (13) is in good agreement with those obtained from continuum elasticity. Notably, we observe variations around the dislocation cores along the $l_{1}$ and $l_{3}$ directions, which reflects the expected strain field behavior in the presence of dislocations. This demonstrates the effectiveness of the Fourier pseudospectral method in solving the APFC model (13), yielding results that closely align with elasticity theory.
@@
Figure 6: Triangular lattice: Strain fields compared with CE along different lines. The red solid lines represent the results from CE theory, while the black circles represent the results from the APFC model (13).
Far-field decay.
To verify the accuracy of the Fourier pseudospectral method for solving the APFC model (13), we consider a larger computational domain with a side length of $640\pi$. We focus on the case of $\varepsilon_{xx}$ in the direction $l_{1}$, noting that the strain field in other directions yields similar results. We demonstrate the results by comparing the decay of $\varepsilon_{xx}$ as a function of the distance $r$ from the dislocation core. As shown in Figure 7, both $\varepsilon_{xx}(\text{APFC})$ and $\varepsilon_{xx}(\text{CE})$, representing the results obtained using the APFC model (13) and continuum elasticity (CE) theory, respectively, decay like $r^{-1}$. The error between them decays as $r^{-2}$, which is consistent with the results shown in 2019Closing .
Figure 7: Triangular lattice: The decay of the strain field $\varepsilon_{xx}$ along the $l_{1}$ direction as a function of the distance $r$ from the dislocation core. The red line represents the results obtained from CE theory, while the green line shows the results from the APFC model (13). The blue line represents the error between them.
@@ 4.2 Three-dimensional BCC lattice
For the three-dimensional BCC lattice described in Example 1 , we first let $\phi_{j}=\phi_{0}$ represent the amplitude of the equilibrium crystal. By minimizing the energy functional given in (15), we obtain $\phi_{0}$ as follows:
$\phi_{0}=\frac{2\gamma+\sqrt{4\gamma^{2}-45vB_{0}}}{45v}.$
@@
A cubic computational domain $L_{x}\times L_{y}\times L_{z}$, with edge lengths $L_{x}=L_{y}=L_{z}=60\pi$, is utilized for numerical simulations. To simulate crystal dislocation, shear stress is applied as in the two-dimensional case. Only horizontal shear stress parallel to the $xy$ plane is applied, with no force exerted in the $z$ direction. Based on $\eta_{j}=\phi_{0}e^{\imath\bm{k}_{j}\cdot\bm{u}(\bm{r})}$, the initial condition $\eta_{j,0}$ of the amplitude functions can be determined by defining $\bm{u}(\bm{r})$ as a displacement to the lattice. Given $\bm{r}:=x\widehat{\bm{x}}+y\widehat{\bm{y}}+z\widehat{\bm{z}}$ representing the spatial position vector, the definition of $\bm{u}(\bm{r})$ is then given by
$\bm{u}(\bm{r})=\left\{\begin{aligned} -u_{x}\widehat{\bm{x}}-u_{y}\widehat{\bm{y}}&,&\frac{L_{z}}{4}<z<\frac{3L_{z}}{4}\\ +u_{x}\widehat{\bm{x}}+u_{y}\widehat{\bm{y}}&,&{\rm elsewhere}\end{aligned}\right$., (42)
where $u_{x}=(a_{x}/L_{x})x$, $u_{y}=(a_{y}/L_{y})y$ and $u_{z}=0$ with $a_{x}=a_{y}=2\pi\sqrt{2}$. With this choice, a strain of
$\varepsilon_{x}=\pm a_{x}/L_{x},\quad\varepsilon_{y}=\pm a_{y}/L_{y},$ (43)
@@
Similar to the two-dimensional case, the numerical accuracy for the BCC lattice is evaluated using the same Fourier pseudospectral setting ($\tau=1\times 10^{-4}$, $h=\pi/128$, $T=6$). The reference solution $\eta_{j}^{*}(\bm{r},T)$ and the numerical error $e_{j,N}$ are defined analogously to (30) and (31) with $j=1,\ldots,6$ and $d=3$. As shown in Figure 8, the errors $e_{j,N}$ for the BCC lattice display spectral convergence with respect to $N$, in agreement with the theoretical results in Section 3.
Figure 8: BCC lattice: numerical error $e_{j,N}$ for $j=1,\ldots,6$ as a function of $N$.